OpenStudy (anonymous):

the coordinates of the of the vertices of a triangle are A(1,3), B(9,10), and C(11,18). find the perimeter of triangle ABC. Find the perimeter of the triangle when the coordinates are multiplied by 3.

OpenStudy (anonymous):

anyone please help meh?

OpenStudy (sohailiftikhar):

do you know the formula of perimeter of triangle ?

OpenStudy (anonymous):

you um add the three sides up?

OpenStudy (sohailiftikhar):

wait

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

have you finally found an answer? it's 2 here :O

OpenStudy (anonymous):

yayyyyyyyyyyyyyy

OpenStudy (sohailiftikhar):

first answer is 36.89

OpenStudy (anonymous):

at least give me the three numbers to add...

OpenStudy (sohailiftikhar):

ok

OpenStudy (sohailiftikhar):

A(1,3), B(9,10), and C(11,18). then the sides will be AB,BC,CA

OpenStudy (sohailiftikhar):

so now you can find the sides ig

OpenStudy (anonymous):

just need the three sides' number that all

OpenStudy (sohailiftikhar):

AB=B-A=(9-1)(10-7)=8i+3j now find it's magnitude

OpenStudy (anonymous):

btw when the sides all multiply by 3 the perometer just get multiplied by 3 right?

OpenStudy (anonymous):

does it have to be that complicated?

OpenStudy (sohailiftikhar):

yes

OpenStudy (sohailiftikhar):

first you have to find sides by distance formula then have to add then that's all .

OpenStudy (anonymous):

\[\sqrt{(y_1+y_2)^2+(x_1+x_2)^2}\]

OpenStudy (anonymous):

that's the formula for finding a side i THINK

OpenStudy (sohailiftikhar):

yes

OpenStudy (anonymous):

but when i plugged it in i got \(\sqrt{113}\)

OpenStudy (anonymous):

but how do u add them TOGETHER :O

OpenStudy (anonymous):

well... you are a HUMAN calculator, right?

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

Use matrix method

OpenStudy (anonymous):

?

OpenStudy (anonymous):

here's what i need: i need the three numbers for the sides of the triangle, the answers for the 3 questions, and a explanation for the last question.

OpenStudy (anonymous):

Well let me think 2 mins plz

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the formula to find the sides are -- Distance from A to B =  In that way you can find all the sides and then for the parameter the formula is , parameter = AB+BC+AC